# AM-RADIO: Agglomerative Model – Reduce All Domains Into One If you can’t beat them, join them (and then beat them)

### v3

@@ Abstract
A handful of visual foundation models (VFMs) have recently emerged as the backbones for numerous downstream tasks. VFMs like CLIP, DINOv2, SAM are trained with distinct objectives, exhibiting unique characteristics for various downstream tasks. We find that despite their conceptual differences, these models can be effectively merged into a unified model through multi-teacher distillation. We name this approach AM-RADIO (Agglomerative Model – Reduce All Domains Into One). This integrative approach not only surpasses the performance of individual teacher models but also amalgamates their distinctive features, such as zero-shot vision-language comprehension, detailed pixel-level understanding, and open vocabulary segmentation capabilities. In pursuit of the most hardware-efficient backbone, we evaluated numerous architectures in our multi-teacher distillation pipeline using the same training recipe. This led to the development of a novel architecture (E-RADIO) that exceeds the performance of its predecessors and is at least 7x faster than the teacher models. Our comprehensive benchmarking process covers downstream tasks including ImageNet classification, ADE20k semantic segmentation, COCO object detection and LLaVa-1.5 framework.
@@ Results
Figure 1: AM-RADIO - is a multi-teacher distillation framework that efficiently trains new vision foundation models of arbitrary architecture. It unifies unique attributes (like zero-shot text grounding, dense correspondence) of each teacher into a single model that even outperforms them on majority of the tasks.
@@ 1 Introduction
Knowledge Distillation [23] has been a very successful and popular technique for transferring the knowledge of a “teacher” model into a typically smaller “student” model. In the original formulation, both the student and the teacher operate on the same in-domain dataset, and the student simultaneously matches the logits of the teacher, and the GT labels. Instead of using labeled images, an alternative approach is to train the student model to match the features of the teacher model [48, 25, 1, 22, 65, 51, 56].
@@
Alternative to using a smaller student model, [58] employ an iterative learning procedure with a high-capacity model where the “student” of equal or greater capacity than the teacher is trained with heavy augmentation applied to the student. Once trained, they expand the dataset by pseudo-labeling new data using the trained student. They then make the student become the teacher, and repeat the process. An important finding in this work is that the student is capable of surpassing the performance of the teacher.
@@
Again in [23], the authors explore the concept of ensemble distillation, where there are multiple teachers, each of which having restricted domain knowledge. [69] provides an overview of multi-teacher distillation, and proposes that instead of matching the summary of an ensemble of teachers, the student can match the features of each individual teacher via some learned non-shared mapping from the representation space of the student to each teacher. Of interest in their approach is that the student and teacher don’t need to share the same architecture, and also that treating teachers individually yields improved performance.
@@
Recently, the concept of Foundation Models (FMs) [3] has emerged, with the general understanding that these models are large, accurate, and typically very expensive to train. They are meant to represent models that, through training on very large datasets, are broadly applicable to numerous downstream tasks. A seminal example of such models is CLIP [46], which trains on web-scale weakly supervised (image, caption) pairs, and results in exceptional zero-shot performances on a wide array of computer vision benchmarks, such as 76.2% zero-shot top-1 accuracy on ImageNet-1K. While CLIP is firmly a FM, another model, DINOv2 [43] has emerged with broad capabilities, often surpassing CLIP. In particular, for tasks that involve holistic representations for an entire image (e.g., classification), DINOv2 isn’t a clear victor compared to CLIP. However, if we turn our attention to dense tasks (e.g., semantic segmentation), DINOv2 shows clear superiority to CLIP on benchmarks such as ADE20k [68] and Pascal VOC [15]. Separately, SAM (Segment Anything) [31] is gaining popularity for its excellent open-vocabulary semantic segmentation abilities, which we hypothesize means it has strong dense feature representations.
@@
We introduce AM-RADIO (Agglomerative Model – Reduce All Domains Into One) with the goal of learning from multiple foundational models simultaneously. We observe that, when given a student model of sufficient capacity, it is often able to exceed any of its teachers on important metrics such as K-NN ImageNet Top-1 accuracy [57, 8, 43] (refer to [57] for implementation details). In addition to performing well on representative foundational benchmarks, by virtue of the training framework, our student models are able to mimic their teacher models, and thus are able to perform downstream tasks that are otherwise performed by the teachers. Examples of this include CLIP-ZeroShot applications, since the language model trained by CLIP is compatible with our student, and also Segment-Anything tasks, as the student is able to replace the vision encoder and interface with the already-trained mask decoders.
@@
In addition to the student itself being a large (and slow) model, we study the effect of using a more hardware-efficient model architecture. Most works are not directly comparable as they use different training recipes, even when evaluated on the same dataset such as ImageNet-1k. To this end, we evaluate more than 10 promising architectures under the same training recipe for a direct comparison. We reveal that CNN-like architectures are faster but less expressive than ViTs. This has led to the development of a novel hybrid architecture, E-RADIO, that exceeds the performance of its predecessors and is at least 7x faster than teacher models.
@@ 2 Related Work
Knowledge Distillation
The underpinning of our work is based on the method of Knowledge Distillation [23, 30, 4, 42, 5] which aims to train a “student” model using soft targets produced by an already-trained “teacher” model, using the the teacher’s output logits as “soft” labels. Alternatively, distillation can be performed using intermediate network activations [48, 25, 1, 22, 65, 51, 56]. In general, due to the heterogeneous nature of the different teacher foundation models that we employ, we ignore any potential labels coming from the data, and we ignore the logits of teachers, and simply opt to match the feature representations of the teachers before any task-specific processing stages.
Multi-Teacher Distillation
There is also a body of work that studies distilling a student model jointly from multiple teacher models simultaneously [23, 37, 69, 64, 66, 61, 45, 62, 32, 2, 17]. Because of the heterogeneous domains that our teacher models cover, we don’t apply approaches that marginalize teachers into a unified label, and instead map students to each teacher independently using teacher-specific projection heads from the unified student representation. Although the reason behind this method in [69] is different, we find the same overall strategy to be effective. While [56] doesn’t study matching the features of multiple teachers simultaneously, we are able to extend their paradigm via the different projection heads. To preserve drop-in compatibility with teacher frameworks, we eliminate the feature normalization in the loss function.
Distilling Foundation Models
Foundation Models [3] are meant to be generalist models that are trained on massive amounts of data, and are typically resource intensive to train from scratch. In the vein of single-teacher distillation, [43] employ self-distillation to train their smaller variants from the larger teacher. [56] distills their model from a CLIP [46] teacher. Instead of focusing our energy on one teacher in particular, we instead grab high-quality versions of the popular CLIP [46] (using OpenCLIP [27]) model, DINOv2 [43], and SAM [31]. Concurrently with our work, [55] describe a methodology for merging a CLIP model into a pretrained SAM model via distillation, which is, in spirit, quite similar to this approach. In contrast to theirs, we include DINOv2 and also simplify the objective to straight-forward feature matching. Since we don’t rely on the student model to be pre-trained, it also gives us the flexibility to have the student be an architecture distinct from any teacher.
@@ 3 Knowledge Agglomeration
We propose a framework to train a vision foundation model from scratch via multi-teacher distillation as shown in Figure 1. We assume that each teacher brings unique properties to the foundational vision model, and the resulting trained model will integrate these attributes.
@@ 3.1 Overview
As an initial assumption, we expect that the teacher models are reasonably capable of representing a broad swath of images found on the internet, coming from datasets such as ImageNet (1k or 21k) [13], LAION-400M [49] or DataComp-1B [18]. With this in mind, we choose to study 3 seminal teacher models: CLIP [46], DINOv2 [43], and SAM [31] as they have demonstrated outstanding performance over a broad range of tasks (as in CLIP), or specifically strong performance on downstream dense tasks, such as semantic segmentation under linear probe (as in DINOv2), or open-vocabulary segmentation (as in SAM). Because these teacher models come from such diverse domains, we omit any form of supplemental ground truth guidance and treat the aforementioned datasets simply as sources of images. To assess the quality of our models, we adopt a set of representative metrics across a few broad domains.
  • Image level reasoning: (i) k-NN Top-1 accuracy on ImageNet-1K, and (ii) Zero-Shot accuracy using the CLIP teacher’s language model. ”Zero-Shot” refers to CLIP-style classification using prompt engineering with the text encoder (not available for DINOv2). k-NN [57, 8, 43] embeds the model’s summary feature vector for every image in the training set, and then for each validation image, it uses a weighted sum of the $k$ nearest training vectors to elect a label.
@@
  • Large Vision-Language Models: we plug our frozen vision encoder model into LLaVA [36] and evaluate it on a wide set of tasks including GQA [26], TextVQA [50], ScienceQA [41] and VQAv2 [20]. Details in Section 5.6.
@@ 3.2 Distillation Dataset Choice
Config Dataset k-NN Zero Shot ADE20K
A Teacher Best 83.41 77.19 47.53
ImageNet 1K 84.79 77.93 48.00
ImageNet 21K 84.03 77.25 48.70
LAION-400M 82.84 74.85 48.41
DataComp-1B 83.17 76.90 48.42
B Teacher Best 83.41 80.51 47.53
@@
Table 1: Ablation study on the choice of different training datasets. Configuration A: we use DINOv2 ViT-G/14 and OpenCLIP ViT-H/14 teachers and a ViT-L/14 [14] student model with CPE [29]; B: we replace the CLIP teacher with MetaCLIP ViT-H/14 and the student model with ViT-H/14. Both “k-NN” and “Zero Shot” are for ImageNet-1k. ADE20k refers to mIOU linear probe on ADE20k. We also report the maximum metrics coming from either of the teacher models.
@@ 3.3 Loss Formulation
Because we don’t have (broadly applicable) ground truth data for each teacher for each image, we instead opt to match the features coming from each teacher’s vision encoder. In particular, we distinguish between the summary feature vector and the spatial feature vectors for each teacher. Because all of our teachers use the ViT architecture, the spatial features are the feature vectors for each input patch coming from the final layer of the model. The summary feature is computed differently based on the model. For CLIP (we use OpenCLIP [27]) and DINOv2, we use the “class token” as the summary feature vector, and for SAM we use the average over the spatial features.
@@
$L_{cos}(x,y)$ $=1-\frac{x^{\intercal}y}{\left|x\right|\left|y\right|}$ (2)
$L_{summary}(x)$ $=\sum_{i}\lambda_{i}L_{cos}(y_{i}^{(s)},z_{i}^{(s)})$
We found empirically that cosine distance loss produced better models compared to L1, MSE, Smooth-L1 [19].
We found that additionally supervising the spatial features of the model by matching the teacher was not only important for downstream dense tasks, it also improved the holistic quality of our model.
@@
Table 2: Ablation over which teachers we supervise the spatial features. We use a ViT-L/14 student model and train on the LAION-400M dataset. Adding this loss term is always beneficial. Dino appears to provide better spatial features than CLIP, but training the student to match both teachers produces the best results.
For matching the spatial features, we employ a combination of cosine similarity and smooth L1. Similar to equation (2) where we found that cosine similarity produced the best results, we found the same to be true for the spatial features, however, since we want to allow our student model to be a drop-in replacement in the teacher frameworks, it’s also important that we match the magnitude of the teacher vectors, and so we introduce smooth L1. In (3) we show the formulation of this loss. Let $h_{i}^{(v)}(x^{\prime}|\Theta_{i}^{(v)})$ be the learned student head for matching teacher feature vectors, and corresponding $t_{i}^{(v)}(x|\Phi_{i}^{(v}))$ be the teacher feature vectors, with $x^{\prime}=f(x|\Theta_{0})$, then the spatial feature loss is:
$L_{\text{match}}(x,y)$ $=\alpha L_{cos}(x,y)+\beta L_{smooth_{l}1}(x,y)$ (3)
$L_{\text{features}}(x)$ $=\sum_{i}\gamma_{i}L_{match}\left(h_{i}^{(v)}(x^{\prime}|\Theta_{i}^{(v)}),t_{i}^{(v)}(x|\Phi_{i}^{(v}))\right)$

### v4

# AM-RADIO: Agglomerative Vision Foundation Model Reduce All Domains Into One

@@ Abstract
A handful of visual foundation models (VFMs) have recently emerged as the backbones for numerous downstream tasks. VFMs like CLIP, DINOv2, SAM are trained with distinct objectives, exhibiting unique characteristics for various downstream tasks. We find that despite their conceptual differences, these models can be effectively merged into a unified model through multi-teacher distillation. We name this approach AM-RADIO (Agglomerative Model – Reduce All Domains Into One). This integrative approach not only surpasses the performance of individual teacher models but also amalgamates their distinctive features, such as zero-shot vision-language comprehension, detailed pixel-level understanding, and open vocabulary segmentation capabilities. Additionally, in pursuit of the most hardware-efficient backbone, we evaluated numerous architectures in our multi-teacher distillation pipeline using the same training recipe. This led to the development of a novel architecture (E-RADIO) that exceeds the performance of its predecessors and is at least 6x faster than the teacher models at matched resolution. Our comprehensive benchmarking process covers downstream tasks including ImageNet classification, semantic segmentation linear probing, COCO object detection and integration into LLaVa-1.5.
@@
Figure 1: PCA feature visualization of different models. Our proposed RADIO model can process any resolution and aspect ratio, and produces semantically rich dense encodings. Not only is the model capable of producing state-of-the-art feature representations at arbitrary resolutions, but the adaptor heads may be used to simulate other models at resolutions they don’t support, such as SAM at larger than 1024px.
@@ Results
Figure 2: AM-RADIO - is a multi-teacher distillation framework that efficiently trains new vision foundation models of arbitrary architecture. It unifies unique attributes (like zero-shot text grounding, dense correspondence) of each teacher into a single model that even outperforms them on a majority of the tasks.
@@ 1 Introduction
Model Params Resol- Throughput ImageNet1K Segmentation (linear) Vision-Language (LLaVa-1.5 [37]) SAM [32]
(M) ution Zero-shot k-NN ADE20k VOC GQA POPE TextVQA VQAv2 COCO
OpenCLIP-H/14 [10] 632 224 503 77.19 81.10 40.04 68.03 57.94 83.61 50.48 72.24 -
MetaCLIP-H/14 [61] 632 224 486 80.51 82.12 35.39 62.62 60.57 84.76 53.65 75.71 -
SigLIP-M/14 [69] 428 384 241 82.61 85.16 40.53 70.31 57.70 84.85 56.65 71.94 -
Intern-ViT-6B [9] 5,902 224 63 83.20†† 78.43 47.20 76.85 60.18 84.02 52.45 76.75 -
5,537 448 14 †† 68.64 42.78 74.43 61.19 87.23 60.36 78.83 -
DFN CLIP-H/14 [17] 633 378 170 83.90 85.27 39.00 70.29 61.73 85.91 56.78 78.78 -
OpenAI CLIP-L/14 [48] 305 336 414 75.54 79.80 36.51 67.04 62.20 86.09 57.92 78.49 -
DINOv2-g/14-reg [13] 1,137 224 294 - 83.41 48.68 82.78 61.88 85.62 47.18 76.23 -
SAM-H/16 [32] 637 1024 12 - 22.12 28.08 34.34 49.92 81.76 43.91 57.65 77.18
E-RADIO-L (Ours) 391 512 468 80.73 83.89 48.22 81.64 61.70 85.07 51.47 76.73 76.31
RADIO-ViT-H/16 (Ours) 653 432 158 82.93 86.06 51.34 84.71 63.01 86.20 56.32 79.28 76.23
Table 1: Comparison of vision foundation and RADIO models. “Zero-Shot” and k-NN are computed on ImageNet-1K. ADE20K [72] and VOC (PascalVOC2012) refer to linear probe semantic segmentation mIOU. GQA, POPE (popular), TextVQA, and VQAv2 are obtained via LLaVa 1.5 [37] by replacing the vision encoder. COCO is the instance segmentation metric introduced by [7] to evaluate SAM [32] distillation. RADIO attains the best metrics on most benchmarks, and is competitive with the rest, while E-RADIO enables high quality results in resource constrained settings. Note that Zero-Shot and COCO use teacher’s decoder head that is not finetuned. Throughput computed using NVIDIA A100 GPU, stated resolution, and TensorRT v8601. *Denotes teachers used to train our final RADIO. We failed to export DINOv2-g-reg to TensorRT, so we report DINOv2-g here, which should be fairly close. ††We were unable to get zero shot working using their model code.
@@
Knowledge Distillation [24] has been a very successful and popular technique for transferring the knowledge of a “teacher” model (or ensemble of models) into a typically smaller “student” model. In the original formulation, both the student and the teacher operate on the same in-domain dataset, and the student simultaneously matches the logits of the teacher, and the ground truth labels. Instead of using labeled images, an alternative approach is to train the student model to match the features of the teacher model [50, 26, 1, 23, 67, 53, 58].
@@
Instead of using a smaller student model, [60] employ an iterative learning procedure with a high-capacity model where a student of equal or greater capacity than the teacher is trained with heavy augmentation applied to the student. Once trained, they expand the dataset by pseudo-labeling new data using the trained student. They then make the student become the teacher, and repeat the process. An important finding in this work is that the student is capable of surpassing the performance of the teacher.
@@
The authors of [24] explore the concept of ensemble distillation, where there are multiple teachers, each of which having restricted domain knowledge. [73] provides an overview of multi-teacher distillation, and proposes that instead of matching the summary of an ensemble of teachers, the student can match the features of each individual teacher via some learned non-shared mapping from the representation space of the student to each teacher. Of interest in their approach is that the student and teacher don’t need to share the same architecture, and also that treating teachers individually yields improved performance.
@@
Recently, the concept of Foundation Models (FMs) [3] has emerged, with the general understanding that these models are large, general, and expensive to train. Through training on very large datasets they are broadly applicable to numerous downstream tasks. A seminal example of such models is CLIP [48], which trains on web-scale weakly supervised (image, caption) pairs, and results in exceptional zero-shot performances on a wide array of computer vision benchmarks. While CLIP is firmly a FM, another model, DINOv2 [45] has emerged with broad capabilities, often surpassing CLIP on dense tasks that require strong spatial features, such as ADE20k [72] and Pascal VOC [16]. Separately, SAM (Segment Anything) [32] is gaining popularity for its excellent open-vocabulary instance segmentation abilities, whose vision encoder we hypothesize has strong dense feature representations.
@@
We introduce AM-RADIO with the goal of learning from multiple foundational models simultaneously. We observe that, when given a student model of sufficient capacity, it is often able to exceed any of its teachers on important axes. In addition to performing well on representative foundational benchmarks, by virtue of the training framework, our student models are able to mimic their teacher models, and thus are able to perform downstream tasks that are otherwise performed by the teachers. Examples of this include CLIP-ZeroShot applications, since the language model trained by CLIP is compatible with our student, and also Segment-Anything tasks, as the student is able to replace the vision encoder and interface with the already-trained mask decoders.
@@
We also study the effect of using a more hardware-efficient model architecture. Most works on efficiency are not directly comparable as they use different training recipes, even when evaluated on the same dataset such as ImageNet-1k, and may be over-tuned. To this end, we evaluate more than 10 promising architectures under the same training recipe for a direct comparison. We reveal that CNN-like architectures are faster but struggle to distill ViT VFMs. This led us to the development of a novel hybrid architecture, E-RADIO, that exceeds the performance of its predecessors and is at least 6x faster than teacher models at matched resolution.
@@ 2 Related Work
Knowledge Distillation The underpinning of our work is based on the method of Knowledge Distillation [24, 31, 4, 44, 5] which aims to train a “student” model using soft targets produced by an already-trained “teacher” model, using the the teacher’s output logits as “soft” labels. Alternatively, distillation can be performed using intermediate network activations [50, 26, 1, 23, 67, 53, 58]. In general, due to the heterogeneous nature of the different teacher foundation models that we employ, we ignore any potential labels coming from the data, and we ignore the logits of teachers, and simply opt to match the feature representations of the teachers before any task-specific processing stages.
Multi-Teacher Distillation There is also a body of work that studies distilling a student model jointly from multiple teacher models simultaneously [24, 39, 73, 66, 70, 63, 47, 64, 33, 2, 18]. Because of the heterogeneous domains that our teacher models cover, we don’t apply approaches that marginalize teachers into a unified label, and instead map students to each teacher independently using teacher-specific projection heads from the unified student representation. Although the reason behind this method in [73] is different, we find the same overall strategy to be effective. While [58] doesn’t study matching the features of multiple teachers simultaneously, we are able to extend their paradigm via the different projection heads. To preserve drop-in compatibility with teacher frameworks, we eliminate the feature normalization in the loss function.
Distilling Foundation Models Foundation Models [3] are meant to be generalist models that are trained on massive amounts of data, and are typically resource intensive to train from scratch. In the vein of single-teacher distillation, [45] employ self-distillation to train their smaller variants from the larger teacher. [58] distills their model from a CLIP [48] teacher. Instead of focusing our energy on one teacher in particular, we instead grab high-quality versions of CLIP [48] (using OpenCLIP [28]), DINOv2 [45], and SAM [32]. Concurrently with our work, [57] describe a methodology for merging a CLIP model into a pretrained SAM model via distillation, which is, in spirit, quite similar to our approach. In contrast to theirs, we include DINOv2 and also simplify the objective to straightforward feature matching. Since we don’t rely on the student model to be pre-trained, it also gives us the flexibility to have the student be an architecture distinct from any teacher.
@@ 3 Knowledge Agglomeration
We propose a framework to train a vision foundation model from scratch via multi-teacher distillation as shown in Figure 2. We demonstrate that each teacher brings unique properties to the foundational vision model, and the resulting trained model will agglomerate these attributes.
@@ 3.1 Overview
As an initial assumption, we expect that the teacher models are capable of representing a broad swath of images found on the internet, coming from datasets such as ImageNet (1k or 21k) [14], LAION-400M [51] or DataComp-1B [19]. With this in mind, we choose to study 3 seminal teacher model families: CLIP [48], DINOv2 [45], and SAM [32] as they have demonstrated outstanding performance over a broad range of tasks (as in CLIP), or specifically strong performance on downstream dense tasks, such as semantic segmentation under linear probe (as in DINOv2), or open-vocabulary segmentation (as in SAM). Because these teacher models come from such diverse domains, we omit any form of supplemental ground truth guidance and treat the aforementioned datasets simply as sources of images. To assess the quality of our models, we adopt a set of representative metrics across a few broad domains.
  • Image level reasoning: (i) k-NN Top-1 accuracy on ImageNet-1K, and (ii) Zero-Shot accuracy using the CLIP teacher’s language model [48]. k-NN [59, 8, 45] embeds the model’s summary feature vector for every image in the training set, and then for each validation image, it uses a weighted sum of the $k$ nearest training vectors to elect a label.
@@
  • Large Vision-Language Models: we plug our frozen vision encoder model into LLaVA-1.5 [37] and evaluate it on a wide set of tasks including GQA [27], TextVQA [52], ScienceQA [43] and VQAv2 [21]. Details in Section 5.4.
  • SAM-COCO instance segmentation: From [7], we adopt their COCO instance segmentation methodology to evaluate our ability to replicate SAM visual features.
Results on these tasks, both for teacher models and our AM-RADIO variants, are summarized in Table 1.
3.2 Adaptor Heads
We opt for simplicity in design of the adaptor heads, and leave alternative architectures as future work. To this end, we employ a simple 2-layer MLP, with a LayerNorm and GELU in between. The input dimension is the student embedding dimension, the intermediate dimension is the maximum embedding dimension of all teachers, and the output dimension matches the specific teacher. For each teacher, we employ two heads, one for the summary vector, and one for the spatial features.
@@ 3.3 Distillation Dataset Choice
Dataset k-NN Zero Shot ADE20K
@@
Table 2: Ablation study on the choice of training dataset. We use MetaCLIP ViT-H/14 [15] and DINOv2 ViT-g/14 teachers, and a ViT-L/14 student model with CPE [30]. Both “k-NN” and “Zero Shot” are for ImageNet-1k. ADE20k refers to mIOU linear probe on ADE20k.
@@ 3.4 Loss Formulation
Because we don’t have ground truth data for each teacher for each image, we instead opt to match the features coming from each teacher’s vision encoder. In particular, we distinguish between the summary feature vector and the spatial feature vectors for each teacher. The summary feature is computed differently based on the model. For CLIP and DINOv2, we use the “class token” as the summary feature vector, and we don’t match a summary for SAM.
@@
$L_{\text{summary}}(x)$ $=\sum_{i}\lambda_{i}L_{\text{cos}}(y_{i}^{(s)},z_{i}^{(s)})$ (2)
We found empirically that cosine distance loss produced better models compared to L1, MSE, Smooth-L1 [20]. Additionally, supervising the spatial features of the model by matching the teacher was not only important for downstream dense tasks, but also improved the holistic quality of our model.
@@
Table 3: Ablation over which teachers we supervise the spatial features. We use a ViT-L/14 student model and train on the LAION-400M dataset. Adding this loss term is always beneficial. DINOv2 appears to provide better spatial features than CLIP, but training the student to match both teachers produces the best results. We don’t ablate SAM as we solely want it for its spatial features.
For matching the spatial features, we employ a combination of cosine similarity and smooth L1. Similar to equation (2) where we found that cosine similarity produced the best results, we found the same to be true for the spatial features. However, we want to allow our student model to be a drop-in replacement in the teacher frameworks, thus it’s important that we match the magnitude of the teacher vectors, and so we include smooth L1. In (3) we show the formulation of this loss. Let $h_{i}^{(v)}(x^{\prime}|\Theta_{i}^{(v)})$ be the learned student head for matching teacher feature vectors, and corresponding $t_{i}^{(v)}(x|\Phi_{i}^{(v)})$ be the teacher feature vectors, with $x^{\prime}=f(x|\Theta_{0})$, then the spatial feature loss is:
$L_{\text{match}}(x,y)$ $=\alpha L_{\text{cos}}(x,y)+\beta L_{\text{smooth-l1}}(x,y)$ (3)
$L_{\text{features}}(x)$ $=\sum_{i}\gamma_{i}L_{\text{match}}\left(h_{i}^{(v)}(x^{\prime}|\Theta_{i}^{(v)}),t_{i}^{(v)}(x|\Phi_{i}^{(v}))\right)$